add: (8/(6x+24)) + (1/(30x+180))
Answer by Nihal@SriLanka(21)

I believe what you are expected to do is simplify the expression.
Here's how you proceed.

8/(6x+24)+1/(30x+180)

= 8/6(x+4)+1/30(x+6)

= 4/3(x+4)+1/30(x+6)

= 4*10(x+6) + (x+4)/30(x+4)(x+6)

= 40x + 240 + x + 4/30(x+4)(x+6)

= (41x + 244) /30(x+4)(x+6)

And that's it, no further simplification being possible.